# Evidence for valence-bond pairing in a low-dimensional two-orbital system

### v1

@@ Abstract
Valence bond (VB) states as the formation mechanism of Cooper pairs, eventually leading to high-temperature superconductivity, remain a controversial topic. Although various VB-like states find variational relevance in the description of specific spin models and quantum spin liquids, in the realm of many-body fermionic Hamiltonians, the evidence for such states as ground states wave functions remains elusive, challenging the valence-bond pairing mechanism. Here, we present evidence of a VB ground state with pairing tendencies, particularly at finite doping. We achieved this for the generic two-orbital Hubbard model in low dimension, where the VB states can be associated with the presence of the topological order manifested by edge states. Utilizing density-matrix renormalization group calculations, the study reveals key properties akin to those observed in superconductors’ phase diagrams, such as pairing restricted to the regime of small but nonzero doping, presence of coherent pairs, and density oscillations in the charge sector.
@@
In 1987, just one year after the discovery of high-temperature superconductivity (high-$T_{\mathrm{c}}$ SC) in cuprates [1], Philip W. Anderson proposed [2] his famous resonating valence bond (RVB) state as the ground state wave-function to describes the properties of such compounds. In essence, the RVB represents a quantum liquid of valence bonds, i.e., a collection of spin singlets, $|\phi\rangle=P_{D}\sum_{j,\ell}\left[a(j-\ell)\,c^{\dagger}_{j,\uparrow}c^{\dagger}_{\ell,\downarrow}\right]^{L/2}|0\rangle$, distributed over the lattice in a way that preserves its spatial symmetries (i.e., no long-range order). Here $|0\rangle$ represent the vacuum, $c^{\dagger}_{\ell,\sigma}$ is the creation operator of an electron with spin projection $\sigma$ at site $\ell$, $P_{D}$ is the Gutzwiller projector preventing double occupancy, while $a(r)$ are some (lattice specific) coefficients. Note that the singlets are not necessarily nearest neighbours and can span over a few lattice sites, though their amplitude is expected to decay exponentially with distance. Such a state can, in principle, describe the Mott insulators’ spin arrangement for the half electronic filing [3] and, more importantly, allow for mobile Cooper pairs under hole or electron doping (with each pair of holes/electrons ”replacing” one of the singlets) [4, 5].
@@
Anderson’s idea has heavily influenced the strongly correlated community for almost 40 years. It sparked enormous interest in the field of quantum magnetism, especially in the antiferromagnetic (AFM) Heisenberg models, which properly describe the main experimental findings of undoped cuprates [6, 7] and parent compounds of some iron-based superconductors [8]. However, RVB and even more generic valence bond solids (VBS, which break some of the lattice symmetry) are rare as the ground states wave-functions of the many-body systems. For the spin models, notable examples of such states exist: (i) it is by now established that quantum spin liquids [9] realized in geometrically frustrated magnets [10, 11] can be described by the RVB wave function. (ii) The ground states of the one-dimensional (1D) $S=1/2$ Heisenberg model with nearest- and next-nearest-neighbour interaction (the so-called Majumdar-Ghosh model) [12, 13] or the two-dimensional (2D) Shastry-Sutherland model [14] are an exact VBS states. (iii) Finally, it was shown [15, 16] that the $S=1$ Heisenberg chain with biquadratic interactions can be thought of as a collection of coupled $S=1/2$-like singlets (Fig. 1a). The latter is encapsulated in the Affleck-Kennedy-Lieb-Tasaki state (AKLT state), which hosts the famous topological Haldane edge states. It is important to note that the AKLT state is a perfect VBS (with correlations vanishing at a distance $r\simeq 2$). At the same time, the plain $S=1$ Heisenberg model in 1D resembles a gapped RVB state with exponentially decaying correlations [17].
@@
Figure 1: Valence-bond induced pairing. a Sketch of the AKLT state (top) and the orbital-RVB state of the doped one-dimensional two-orbital Hubbard model (lower). Paired holes are presented as circles. b Interaction $U$ – doping $x$ phase diagram of the string order parameter $\mathcal{O}_{\mathrm{s}}$ evaluated in the bulk (at distance $\ell=L/2$) of the $L=60$ sites system. The points depict values at which extrapolated to $L\to\infty$ spin gap $\Delta_{\mathrm{S}}$ opens or closes. c Phase diagram for binding energy $\epsilon_{\mathrm{B}}=E_{\mathrm{gs}}(N)-2E_{\mathrm{gs}}(N-1)+E_{\mathrm{gs}}(N-2)$, where $E_{\mathrm{gs}}(N)$ is the energy of the fermionic system with $N$ electrons ($L=60$). Points depict values at which the extrapolated to $L\to\infty$ binding energy crosses the zero value. Lines are a guide to the eye.
@@
In the context of fermionic Hamiltonians (with the fermionic Hubbard model as a prime example), evidence of a VB-like state as the ground state of the many-body system is still lacking. Although such states are anticipated to capture many properties of quantum paramagnets (systems with short-ranged AFM correlations), the challenge is demonstrating that the ground state has a form of VB liquid. Consequently, the valence-bond pairing mechanism, as envisioned by Anderson, is questioned. Here, we report the evidence that the ground state of the generic two-orbital Hubbard model in low-dimension realizes a VB-like state. The latter maintains the topological properties of the AKLT state and, in addition, becomes superconducting at finite doping. Namely, we show that the topology and pairing are intertwined in the doped fermionic Haldane chain. Although limited to 1D considerations, our density-matrix renormalization group (DMRG) calculations reveal most of the properties expected of the high-$T_{\mathrm{c}}$ phase diagram, i.e., (i) a large region of finite (negative) binding energy in the interaction-doping phase diagram; (ii) long-range of pair-pair correlations; and (iii) pair-density-wave (PDW). We also show that (iv) all of these phenomena are induced by the VB ground state akin to the AKLT state of $S=1$ Heisenberg chains. Since the latter exhibits topological properties, the presence of VB-type states can be easily identified via the presence of the topological order.
@@
Our investigation is based on the two-orbital ($\gamma=0,1$) Hubbard-Kanamori model on the 1D lattice:
@@
with $P^{\dagger}_{\gamma\ell}=c^{\dagger}_{\gamma\uparrow\ell}c^{\dagger}_{\gamma\downarrow\ell}$. In the following, we will consider its most generic version, with band degeneracy. The first term describes the system’s kinetic energy (with $t=0.5\,\mathrm{[eV]}$ and kinetic energy span $W=4t$ as a unit of energy). The second term describes intra- ($U$) and inter-orbital ($U^{\prime}$) on-site electron repulsion. The last term originates in multi-orbital physics: $J_{\mathrm{H}}$ accounts for the ferromagnetic Hund coupling between spins $\mathbf{S}_{\gamma\ell}$ at different orbitals, maximizing the total on-site spin. The above model preserves SU(2) symmetry (provided that $U^{\prime}=U-5/2J_{\mathrm{H}}$ [18]), and we will consider $J_{\mathrm{H}}/U=1/4$ in the $S^{z}_{\mathrm{tot}}=0$ magnetization sector for various hole doping levels $x=1-\overline{n}$, where $\overline{n}=N/2L$ is electron density (with $N$ as a number of electrons in the $L$ site system). The quasi-1D (ladders) and 2D versions of the above model are extensively used in the context of various correlated superconductors like iron pnictides, chalcogenides, ruthenates, iridates, as well as heavy-fermion materials. The single-band version of the above Hamiltonian is also considered to be a minimal model for cuprates. In the following, we will present results obtained with the help of the DMRG method on the open boundary system (see Supplementary Information for details).
Figure 2: String order correlations $\mathcal{O}_{\mathrm{s}}$. a Distance dependence of $\mathcal{O}_{\mathrm{s}}(\ell)$ for various interaction strength $U/W$ and half-filled system $x=0.0$. Change in the behaviour of $\mathcal{O}_{\mathrm{s}}$ is evident for $U/W\simeq 0.5$, i.e., for the predicted [19] transition to the topological Haldane phase in which the string order is not decaying for $\ell\to\infty$. b,c Similar results for hole doped system: $x=0.1$ and $x=0.2$, respectively. All results evaluated for $L=80$ site system. d,e Finite-size scaling ($L=40,60,80$) for doped systems, d $x=0.1$ and e $x=0.2$, close to the transition. f Amplitude of the charge density oscillations ($L=60$). Lines represent the same guide to the eye as in Fig. 1.
@@
$|1_{i}\rangle=|\uparrow_{i}\uparrow_{i}\rangle\,,\quad|-1_{i}\rangle=|\downarrow_{i}\downarrow_{i}\rangle\,,$
$|0_{i}\rangle=\frac{1}{\sqrt{2}}\left(|\uparrow_{i}\downarrow_{i}\rangle+|\downarrow_{i}\uparrow_{i}\rangle\right)\,,$
which are coupled in a valence bond way between sites, $(|\uparrow_{i}\downarrow_{i+1}\rangle-|\downarrow_{i}\uparrow_{i+1}\rangle)$. Here $|1,0,-1\rangle$ represent $S=1$ states at site $i$, while $|\sigma_{i}\sigma^{\prime}_{j}\rangle$ with $\sigma_{i}=\uparrow,\downarrow$ can be thought as a state of two electrons at different orbitals in the context of the two-orbital Hubbard model. The above VBS (AKLT state) is not an exact ground state of the isotropic $S=1$ AFM model. Still, it can be adiabatically connected to it without closing the spin gap and preserving its unique properties, i.e., the presence of the topologically protected Haldane edge states (unpaired by valence bonds $S=1/2$ states at the boundary of the open system). It was recently shown [19] that such a description of the two-orbital Hubbard model in 1D is valid even for a relatively small value of interaction, $U\simeq W/2$, in the region where magnetic moments are not fully developed $\mathbf{S}^{2}\ll 2$ and charge fluctuations are still present. Such behaviour can be monitored with the help of the string order correlation function:
@@
with $S^{z}_{m}=S^{z}_{0m}+S^{z}_{1m}$ as the total spin at site $m$, which serves as the order parameter of the AKLT state in the $\ell\to\infty$ limit (i.e., breaking of the discrete $Z_{2}\times Z_{2}$ hidden symmetry).
@@
Introduction of the holes into the AKLT state is a nontrivial task. In principle, two scenarios are possible: the formation of ”rigid” on-site holes $|h_{i}h_{i}\rangle$ in the AFM $S=1$ background or the pair of holes transforming two of the $S=1$ on-site triplets into two $S=1/2$ objects, e.g.,
$|\uparrow_{i}\downarrow_{j}\rangle-|\downarrow_{i}\uparrow_{j}\rangle\quad\to\quad|\uparrow_{i}h_{j}\rangle-|h_{i}\downarrow_{j}\rangle\,.$ (3)
@@
In the atomic limit, the former scenario is favoured by the Hund exchange $J_{\mathrm{H}}$ since it’s maximalising the average magnetic moments. On the other hand, the latter case - with one hole per site - is preferred by the inter-orbital repulsion $U^{\prime}$ term in the Hamiltonian. To resolve this issue, we monitor the behaviour of the ground state with doping $x$ and interaction $U$ by evaluating the string order parameter $\mathcal{O}_{s}(\ell)$. Detailed distance dependence is presented in Fig. 2, while Fig. 1b depicts the phase diagram [obtained from the bulk value of $\mathcal{O}_{\mathrm{s}}(L/2)$]. For dopings $x\lesssim 0.35$, one can observe a region where the string order parameter is finite, excluding the ”rigid” hole scenario which would break the AFM chain. More importantly, our results are also consistent with one hole per site scenario, i.e., with (two) holes replacing one of the valence bonds (see Eq. 3). For a few holes away from half-filling, $x=0$, such state was coined the orbital-RVB [20, 21, 22]. Here, we extend this definition to rather large doping levels, $x\simeq 0.35$, and show that its properties are consistent with the valence-bond pairing mechanism.
For $x=0$ (half-filling), one can observe (Fig. 1b) a finite string-order parameter for all $U/W\gtrsim 0.5$ (with $U\to\infty$ limit given by the $S=1$ Heisenberg model), while for $x\neq 0$ this is true only in a finite region of interaction strength $U$. One can understand the lower (at small $U$) topological phase transition (from trivial paramagnetic to topological orbital-RVB state) as an effect of interaction $U$ strengthening the magnetic moments $\mathbf{S}$ and decreasing the charge fluctuations. On the other hand, the upper (at large $U$) transition is associated with the change in the spin-spin correlations from AFM (incommensurate at finite $x$) to ferromagnetic (FM) ordering due to double-exchange like physics in the large Hund limit for $x\neq 0$ [23, 24]. In Supplementary Information, we also present the analysis of the correlations between the system’s edges, confirming the presence of topologically protected edge states, and the spin-spin magnetic structure factor analysis confirming the AFM to FM transition at large $U$.
The spatial dependence of the string order parameter $\mathcal{O}_{\mathrm{s}}(\ell)$ indicates the presence of pronounced oscillations for $x\neq 0$ (see Fig. 2). We associate them with the presence of Friedel oscillations in the charge sector (due to the open boundary system), visible also in the site-resolved electron density $n_{\ell}=n_{0\ell}+n_{1\ell}$ (see the Supplementary Information). In Fig. 2f, we present the spatial standard deviation of the density $\sigma_{n}=\sqrt{(1/L)\sum_{\ell}n_{\ell}^{2}-\overline{n}^{2}}$, related to the amplitude $A$ of the cosine-wave, $\sigma_{n}\simeq A/\sqrt{2}$. In the topologically trivial region, $U\lesssim W$ and $x\gtrsim 0.2$ we find the ”standard” $2k_{\mathrm{F}}=\pi\overline{n}=\pi(1-x)\propto\pi x$ charge-density-wave oscillations of weakly interacting system. Interestingly, in the region where we find finite $\mathcal{O}_{\mathrm{s}}\neq 0$ ($U\sim 2W$, $x\lesssim 0.35$), our results indicate that the charge oscillates with $2\pi x$ wave-vector. The latter is consistent with a pair-density wave (PDW) with $4k_{\mathrm{F}}$ oscillations [25, 26, 27] (leading to two doped holes per one minimum in oscillation). This indicated that the PDW accompanies the doped orbital-RVB and is most pronounced at doping $x\simeq 0.15$.
Our results show that the orbital-RVB state remains robust even for large doping $x\simeq 0.3$ at $U\simeq 4W$. In Fig. 1c, we contrast this behaviour with the binding energy $\epsilon_{\mathrm{B}}$, where a negative value signals the presence of bound pairs of holes (or electrons for $x<0$, due to the particle-hole symmetry of the considered model). The significant overlap in the interaction-doping phase diagram between the finite string order, $\mathcal{O}_{\mathrm{s}}\neq 0$, and negative binding energy, $\epsilon_{\mathrm{B}}<0$, indicates that the VB structure of the ground state and bounded pairs coexist. Consequently, the latter are correlated due to the former (see Fig. 1a). The above behaviour is in accord with the scenario envisioned by Anderson’s proposal: upon doping, the system minimises its energy by breaking a minimal amount of coupled singlets (valence bonds) [28]. Investigating the spin gap $\Delta_{\mathrm{S}}$ can also confirm this picture. In the $U\gg W$ limit and for $x=0$, i.e., when the $S=1$ Heisenberg model should properly describe the low-energy excitations, one expects the famous Haldane gap $\Delta_{\mathrm{S=1}}\simeq 0.41J$ of the AKLT state, with spin exchange $J=2t^{2}/(U+J_{\mathrm{H}})$ [19]. For finite doping, our results indicate (Fig. 3a) that the spin gap remains finite $\Delta_{\mathrm{S}}\neq 0$ in the same region where the string order is finite (see also points in Fig. 1b). Interestingly, the VB-induced pairing (as measured by negative binding energy) is found for the electron-electron interaction strengths $U$, which are ”just below” the large interaction expansion, i.e., below the $J(U)\propto 1/U$ energy scale which gives the proper description of the spin excitations.
@@
Figure 3: Spin gap and binding energy. Extrapolated to $L\to\infty$ a spin gap $\Delta_{\mathrm{S}}=E_{\mathrm{gs}}(S^{z}_{\mathrm{tot}}=2)-E_{\mathrm{gs}}(S^{z}_{\mathrm{tot}}=0)$ and b binding energy $-\epsilon_{\mathrm{B}}$. Dashed line in a indicates Haldane gap $\Delta_{\mathrm{S=1}}\simeq 0.41J$ with $J=2t^{2}/(U+J_{\mathrm{H}})$.

### v2

# Evidence for valence-bond pairing in a one-dimensional two-orbital system

@@ Abstract
Valence bond (VB) states as the formation mechanism of Cooper pairs, eventually leading to high-temperature superconductivity, remain a controversial topic. Although various VB-like states find variational relevance in the description of specific spin models and quantum spin liquids, in the realm of many-body fermionic Hamiltonians, the evidence for such states as ground states wave functions remains elusive, challenging the valence-bond pairing mechanism. Here, we present evidence of a VB ground state with pairing tendencies, particularly at finite doping. We achieved this for the generic two-orbital Hubbard model in low dimension, where the VB states can be associated with the presence of the topological order manifested by edge states.
@@
Introduction.–In 1987, just one year after the discovery of high-temperature superconductivity (high-$T_{\mathrm{c}}$ SC) in cuprates [1], Philip W. Anderson proposed [2] his famous resonating valence bond (RVB) state as the ground state wavefunction to describes the properties of such compounds. In essence, the RVB represents a quantum liquid of valence bonds, i.e., a collection of spin singlets, distributed over the lattice in a way that preserves its spatial symmetries. Note that the singlets are not necessarily nearest neighbours and can span over a few lattice sites, though their amplitude is expected to decay exponentially with distance. Such a state can, in principle, describe the Mott insulators’ spin arrangement for the half electronic filing [3] and, more importantly, allow for mobile Cooper pairs under hole or electron doping (with each pair of holes/electrons ”replacing” one of the singlets) [4, 5].
@@
Anderson’s idea sparked enormous interest in the antiferromagnetic (AFM) Heisenberg models, which properly describe the main experimental findings of undoped cuprates [6, 7] and parent compounds of some iron-based superconductors [8]. However, RVB and even more generic valence bond solids (VBS, which break some of the lattice symmetry) are rare as the ground states wavefunctions of the many-body systems. For the spin models, notable examples of such states exist: (i) it is by now established that quantum spin liquids [9] realized in geometrically frustrated magnets [10, 11] can be described by the RVB wave function. (ii) The ground states of the one-dimensional (1D) $S=1/2$ Heisenberg model with nearest- and next-nearest-neighbour interaction (the so-called Majumdar-Ghosh model) [12, 13] or the two-dimensional (2D) Shastry-Sutherland model [14] are exact VBS states. (iii) Finally, it was shown [15, 16] that the $S=1$ Heisenberg chain with biquadratic interactions can be thought of as a collection of coupled $S=1/2$-like singlets (Fig. 1a). The latter is encapsulated in the Affleck-Kennedy-Lieb-Tasaki state (AKLT state), which hosts the famous topological Haldane edge states. It is important to note that the AKLT state is a perfect VBS. At the same time, the plain $S=1$ Heisenberg model in 1D resembles a gapped RVB state with exponentially decaying correlations [17].
@@
Figure 1: a Sketch of the AKLT state (top) and the orbital-RVB state of the doped one-dimensional two-orbital Hubbard model (lower). Paired holes are presented as circles. b Interaction $U$ – doping $x$ phase diagram of the string order parameter $\mathcal{O}_{\mathrm{s}}$ evaluated in the bulk (at distance $\ell=L/2$) of the $L=60$ sites system. The points depict values at which extrapolated to $L\to\infty$ spin gap $\Delta_{\mathrm{S}}$ opens or closes. c Phase diagram for binding energy $\epsilon_{\mathrm{B}}=E_{\mathrm{gs}}(N)-2E_{\mathrm{gs}}(N-1)+E_{\mathrm{gs}}(N-2)$, where $E_{\mathrm{gs}}(N)$ is the energy of the fermionic system with $N$ electrons ($L=60$). Points depict values at which the extrapolated to $L\to\infty$ binding energy crosses the zero value. Lines are a guide to the eye.
@@
In the context of fermionic Hamiltonians, evidence of a VB-like state as the ground state of the many-body system is still lacking. Although such states are anticipated to capture many properties of quantum paramagnets, the challenge is demonstrating that the ground state has a form of VB liquid. Here, we report the evidence that the ground state of the generic two-orbital Hubbard model in low-dimension realizes a VB-like state. The latter maintains the topological properties of the AKLT state and, in addition, becomes superconducting at finite doping. Namely, we show that the topology and pairing are intertwined in the doped fermionic Haldane chain. Although limited to 1D considerations, our density-matrix renormalization group (DMRG) calculations reveal most of the properties expected of the high-$T_{\mathrm{c}}$ phase diagram, i.e., (i) a large region of finite binding energy in the interaction-doping phase diagram; (ii) long-range of pair-pair correlations; and (iii) pair-density-wave (PDW). We also show that (iv) all of these phenomena are induced by the VB ground state akin to the AKLT state of $S=1$ Heisenberg chains. Since the latter exhibits topological properties, the presence of VB-type states can be easily identified via the presence of the topological order.
@@
Orbital-RVB.–Our investigation is based on the two-orbital ($\gamma=0,1$) Hubbard-Kanamori model on the 1D lattice:
@@
with $P^{\dagger}_{\gamma\ell}=c^{\dagger}_{\gamma\ell\uparrow}c^{\dagger}_{\gamma\ell\downarrow}$. In the following, we will consider its most generic version, with band degeneracy. The first term describes the system’s kinetic energy (with kinetic energy span given by $W=4t$). The second term describes intra- ($U$) and inter-orbital ($U^{\prime}$) on-site electron repulsion. The last term originates in multi-orbital physics: $J_{\mathrm{H}}$ accounts for the ferromagnetic Hund coupling between spins $\mathbf{S}_{\gamma\ell}$ at different orbitals, maximizing the total on-site spin. The above model preserves SU(2) symmetry (provided that $U^{\prime}=U-5/2J_{\mathrm{H}}$ [18]), and we will consider $J_{\mathrm{H}}/U=1/4$ in the $S^{z}_{\mathrm{tot}}=0$ magnetization sector for various hole doping levels $x=1-\overline{n}$, where $\overline{n}=N/2L$ is electron density (with $N$ as the number of electrons in a $L$ site system). The quasi-1D (ladders) and 2D versions of the above model are extensively used in the context of various correlated superconductors like iron pnictides, chalcogenides, ruthenates, iridates, as well as heavy-fermion materials. In the following, we will present results obtained with the help of the DMRG method on the open boundary system (see Supplemental Material [19] and references [20, 21, 22] therein for details).
@@
$|1_{i}\rangle=|\uparrow_{0i}\uparrow_{1i}\rangle\,,\quad|-1_{i}\rangle=|\downarrow_{0i}\downarrow_{1i}\rangle\,,$
$|0_{i}\rangle=\frac{1}{\sqrt{2}}\left(|\uparrow_{0i}\downarrow_{1i}\rangle+|\downarrow_{0i}\uparrow_{1i}\rangle\right)\,,$
which are coupled in a valence bond way between sites, $(1/\sqrt{2})(|\uparrow_{0i}\downarrow_{1i+1}\rangle-|\downarrow_{0i}\uparrow_{1i+1}\rangle)$. Here $|1,0,-1\rangle$ represent $S=1$ states at site $i$, while $|\sigma_{\gamma i}\sigma_{\gamma^{\prime}j}\rangle$ with $\sigma_{i}=\uparrow,\downarrow$ can be thought as a spin configuration of two electrons at different orbitals $\gamma\,,\gamma^{\prime}$ in the context of the two-orbital Hubbard model. The above VBS (AKLT state) is not an exact ground state of the isotropic $S=1$ AFM model. Still, it can be adiabatically connected to it without closing the spin gap and preserving its unique properties, i.e., the presence of the topologically protected Haldane edge states (unpaired by valence bonds $S=1/2$ states at the boundary of the open system). It was recently shown [23] that such a description of the two-orbital Hubbard model in 1D is valid even for a relatively small value of interaction, $U\simeq W/2$, in the region where magnetic moments are not fully developed $\mathbf{S}^{2}\ll 2$ and charge fluctuations are still present. Such behaviour can be monitored with the help of the string order correlation function:
@@
with $|\mathrm{gs}\rangle$ as the ground state wavefunction and $S^{z}_{m}=S^{z}_{0m}+S^{z}_{1m}$ as the total spin at site $m$, which serves as the order parameter of the AKLT state in the $\ell\to\infty$ limit (i.e., breaking of the discrete $Z_{2}\times Z_{2}$ hidden symmetry).
Figure 2: a Distance dependence of string order correlations function $\mathcal{O}_{\mathrm{s}}(\ell)$ for various interaction strength $U/W$ and $x=0.1$ hole doped system ($L=80$ data). b Amplitude of the charge density oscillations ($L=60$). Lines represent the same guide to the eye as in Fig. 1.
@@
Introduction of the holes into the AKLT state is a nontrivial task. In principle, two scenarios are possible in the atomic limit: the formation of ”rigid” on-site holes $|h_{0i}h_{1i}\rangle$ in the AFM $S=1$ background or the pair of holes transforming two of the $S=1$ on-site triplets into two $S=1/2$ objects, e.g.,
$|\uparrow_{0i}\downarrow_{1i}\rangle-|\downarrow_{0i}\uparrow_{1i}\rangle\quad\to\quad|\uparrow_{0i}h_{1i}\rangle-|h_{0i}\uparrow_{1i}\rangle.$
@@
In the atomic limit $t=0$, the former scenario is favoured by the Hund exchange $J_{\mathrm{H}}$ since it’s maximizing the average magnetic moments. On the other hand, the latter case - with one hole per site - is preferred by the inter-orbital repulsion $U^{\prime}$ term in the Hamiltonian. To resolve this state in the many-body system $t\neq 0$, we monitor the behaviour of the ground state with doping $x$ and interaction $U$ by evaluating the string order parameter $\mathcal{O}_{s}(\ell)$. Detailed distance dependence is presented in Fig. 2a, while Fig. 1b depicts the phase diagram obtained from the bulk value of $\mathcal{O}_{\mathrm{s}}(L/2)$ (see also Supplemental Material [19] for additional results). For dopings $x\lesssim 0.35$, one can observe a region where the string order parameter is finite, excluding two holes at the same site (but different orbitals $|h_{i0}h_{i1}\rangle$), which would break the AFM chain into two parts. More importantly, our results are consistent with a one hole per site scenario, i.e., with (two) holes replacing one of the valence bonds. For a few holes away from half-filling, $x=0$, such a state was coined the orbital-RVB [24, 25, 26]. Here, we extend this definition to rather large doping levels, $x\simeq 0.35$, and show that its properties are consistent with the valence-bond pairing mechanism.
@@
Figure 3: Extrapolated to $L\to\infty$ a spin gap $\Delta_{\mathrm{S}}=E_{\mathrm{gs}}(S^{z}_{\mathrm{tot}}=2)-E_{\mathrm{gs}}(S^{z}_{\mathrm{tot}}=0)$ and b binding energy $-\epsilon_{\mathrm{B}}$. Dashed line in a indicates Haldane gap $\Delta_{\mathrm{S=1}}=0.41J$ with $J=2t^{2}/(U+J_{\mathrm{H}})$.
For $x=0$ (half-filling), one can observe (Fig. 1b) a finite string-order parameter for all $U/W\gtrsim 0.5$ (with $U\to\infty$ limit given by the $S=1$ Heisenberg model), while for $x\neq 0$ this is true only in a finite region of interaction strength $U$. One can understand the lower (at small $U$) topological phase transition (from trivial paramagnetic to topological orbital-RVB state) as an effect of interaction $U$ strengthening the magnetic moments $\mathbf{S}$ and decreasing the charge fluctuations. On the other hand, the upper (at large $U$) transition is associated with the change in the spin-spin correlations from AFM (incommensurate at finite $x$) to ferromagnetic (FM) ordering due to double-exchange like physics in the large Hund limit for $x\neq 0$ [27, 28]. In Supplementary Material (see also references [29, 30, 31, 32, 33, 34, 35, 36, 37] therein), we also present the analysis of the correlations between the system’s edges, confirming the presence of topologically protected edge states, and the spin-spin magnetic structure factor analysis confirming the AFM to FM transition at large $U$.
Since the orbital-RVB state shares its topological properties with the ”standard” AKLT state, one expects that finite string order $\mathcal{O}_{s}\neq 0$ is accompanied by the presence of the reminiscence of the Haldane gap [$\Delta_{\mathrm{S=1}}\simeq 0.41J$ for the $x=0$ AKLT state, with spin exchange $J=2t^{2}/(U+J_{\mathrm{H}})$ [23]]. In Fig. 3a, we present the interaction $U$ dependence of the (extrapolated to thermodynamic limit [19]) spin gap $\Delta_{\mathrm{S}}$. Indeed, our results indicate that $\Delta_{\mathrm{S}}$ remains open in the same region where the string order is finite, also for nonzero doping $x\neq 0$ (see points presented in Fig. 1b, which depict opening and closing of the spin gap).
Figure 4: a Doping $x$ dependence of the pair-pair correlation function $C_{\mathrm{P}}(\ell)$ for $U/W=1.3$ ($L=80$ data). b Finite-size scaling ($L=40,60,80$) of pair-pair correlation length $\xi_{\mathrm{P}}$ obtained from $C_{\mathrm{P}}(\ell)\propto\exp(-\ell/\xi_{\mathrm{P}})$ fits (see Supplementary Material for more data [19]). c Phase diagram of the correlation length $\xi_{\mathrm{P}}$ obtained from the fits to the $L=60$ sites system. Points depict parameters presented in panel b. Lines represent the same lines represent guides to the eye as in Fig. 1.
Finally, the spatial dependence of the string order parameter $\mathcal{O}_{\mathrm{s}}(\ell)$ indicates the presence of pronounced oscillations for $x\neq 0$. We associate them with the presence of Friedel oscillations in the charge sector (due to the open boundary system), visible also in the site-resolved electron density $n_{\ell}=n_{0\ell}+n_{1\ell}$ (see the Supplementary Material [19]). In Fig. 2b, we present the spatial standard deviation of the density $\sigma_{n}=\sqrt{(1/L)\sum_{\ell}n_{\ell}^{2}-\overline{n}^{2}}$, related to the amplitude $A$ of the cosine-wave, $\sigma_{n}\simeq A/\sqrt{2}$. In the topologically trivial region, $U\lesssim W$ and $x\gtrsim 0.2$ we find the ”standard” $2k_{\mathrm{F}}=\pi\overline{n}=\pi(1-x)\propto\pi x$ charge-density-wave oscillations of a weakly interacting system. Interestingly, in the region where we find finite $\mathcal{O}_{\mathrm{s}}\neq 0$ ($U\sim 2W$, $x\lesssim 0.35$), our results indicate that the charge oscillates with $2\pi x$ wave-vector. The latter is consistent with a pair-density wave (PDW) with $4k_{\mathrm{F}}$ oscillations [38, 39, 40] (leading to two doped holes per one minimum in oscillation). This indicated that the PDW accompanies the doped orbital-RVB and is most pronounced at doping $x\simeq 0.15$.
Binding energy.–Our results show that the orbital-RVB state remains robust even for large doping $x\simeq 0.3$ at $U\simeq 4W$. In Fig. 1c, we contrast this behaviour with the binding energy $\epsilon_{\mathrm{B}}$, where a negative value signals the presence of bound pairs of holes (or electrons for $x<0$, due to the particle-hole symmetry of the considered model). The significant overlap in the interaction-doping phase diagram between the finite string order, $\mathcal{O}_{\mathrm{s}}\neq 0$, and negative binding energy, $\epsilon_{\mathrm{B}}<0$, indicates that the VB structure of the ground state and bounded pairs coexist. Consequently, the latter are correlated due to the former (see Fig. 1a). The above behaviour is in accord with the scenario envisioned by Anderson’s proposal: upon doping, the system minimises its energy by breaking a minimal amount of coupled singlets (valence bonds) [41]. Interestingly, the VB-induced pairing (as measured by negative binding energy) is found for the electron-electron interaction strengths $U$, which are ”just below” the large interaction expansion, i.e., below the $J(U)\propto 1/U$ energy scale, which gives the proper description of the spin excitations.